OpenStudy (anonymous):

Which function represents a translation of the graph of y=x^2 by 8 units to the right? Which function represents a translation of the graph of y=x^2 by 8 units to the right? @Mathematics

OpenStudy (anonymous):

y=(x^2-8)

OpenStudy (anonymous):

a. y(x+8)^2 b. y=(x-8)^2 c. y=x^2+8 d. y=8x^2 so is it b?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

iam 100%sure